OpenStudy (anonymous):

Find the zeros of g(x)=5(x+6)^2(x^2-36)(x-6)? I'll give a metal to whoever can solve this for me!

OpenStudy (anonymous):

the equal is supposed to be a plus before the six my bad!

OpenStudy (anonymous):

x=-6 with multiplicity: 2, 6,-6,6 I think

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

To find the roots/zeros of the function, set the function equal to 0 and solve. If any individual factor on the left-hand side of the equation is equal to 0, the entire expression will be equal to 0. Set the first factor equal to 0 and solve. Then, set each of the factors equal to 0 and solve. e.g. (x+6) x+6=0 x=-6 don't forget to take both positive and negative square root possibilities as zeroes.

hero (hero):

Set g(x) = 0 0=5(x+6)^2(x^2-36)(x-6) Divide both sides by 5 0 = (x + 6)^2(x^2 - 36)(x - 6) Use zero product property at this point: 0 = x + 6 0 = x - 6